simplify the expression. 16 + (x + 4)
(a) x + 12
(b) x + 20
(c) x + 4
(d) 4x + 16
simplify the expression. (8 * b) * 4
(a) b + 32
(b) 12b
(c) 32b
(d) 8b + 4
translate the verbal sentence into an equation and solve. the sum of a number and 16 is 31.
(a) 16 + 31 = x,x = 47
(b) x + 16 = 31,x = 15
(c) x + 16 = 31,x = 47
(d) 16 + 31 = x,x = 15
translate the verbal sentence into an equation and solve. the difference of 22 and a number is 8.
(a) y - 22 = 8,y = 30
(b) 22 - y = 8,y = 30
(c) y - 22 = 8,y = 14
(d) 22 - y = 8,y = 14

Answer by ewatrrr(24785)


Hi
simplify the expression. 16 + (x + 4) = x + 20
simplify the expression. (8 * b) * 4 = 32b
the sum of a number and 16 is 31.
x + 16 = 31 |subtracting 16 from both sides of equation
x = 15 | x is equal to 15
the difference of 22 and a number is 8.
22 - y = 8 |subtracting 8 and adding y to both sides of equation
14 = y | y is equal to 14
Note: when solving, we must isolate the variable on one side of the equation,
so as, we can 'read' the solution